Jessica invested £14000 for 5 years at a compound interest rate of 9% per year. How much more could she have earned if the interest was compounded monthly ?
Found 2 solutions by stanbon, TimothyLamb:
Answer by stanbon(75887)

Jessica invested £14000 for 5 years at a compound interest rate of 9% p.a.
How much more could she have earned if the interest was monthly ?
--------------
Yearly Compounding :: 14000(1+(0.09/1))^(1*5) = 21540.74
----
Monthly Compounding :: 14000(1+(0.09/12))^(12*5) = 21919.53
--------
Ans: 21919.53-21540.74 = 378.79
---------------
Cheers,
Stan H.
---------------

Answer by TimothyLamb(4379)

compound-interest formula:
v = p( 1 + (r/n) )^(nt)
where:
v = final value
p = initial (principal) value
r = interest rate (as a decimal)
n = number of compounding periods per year
t = number of years
---
yearly compounding:
a = 14000( 1 + (0.09/1) )^(1*5)
a = 21540.74
---
monthly compounding:
b = 14000( 1 + (0.09/12) )^(12*5)
b = 21919.53
---
b - a = 21919.53 - 21540.74
b - a = 378.79
---
answer:
she could have earned 378.79 more with monthly compounding
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---